The measure of the angle of a triangle are 3 consecutive even integers. Find the measure of each angle. Recall that the sum of the angles in any triangles is 180 degrees
Answer by checkley79(3341)

Let x, x+2 & x+4 be thw three angles.
x+(x+2)+(x+4)=180
3x+6=180
3x=180-6
3x=174
x=174/3
x=58 ans. for the smallest angle.
58+2=60 ans. for the second angle.
58+4=62 ans for the largest angle.
Proof:
58+60+62=180
180=180